• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name cusanno

    Log entry time 08:07:11 on June23,2007

    Entry number 206377

    keyword=Shift Summary

    00:00 Shift starts in the middle of run # 1464, radiator in.

    00:33 Moved radiator OUT

    00:44 We noted NMR on Left Arm pretty noisy, furthermore Left VDC is not very good

    Working on Left VDC problems, it is not due to the gas system, we rebooted fastbus crates, no helps

    Alexandre Camsonne come in to look at the problem

    Asked for Controlled Access

    03:45 Alexandre Camsonne and Xiaodong Jiang in Controlled Access

    04:15 Problem fixed

    04:25 Data taking restored

    05:10 during the checklist, we found FPP threshold at zero. So we also moved radiator IN to keep the 4:1 ratio between IN and OUT

    07:25 Beam trip due to IOC failure

    07:32 Beam back

    08:00 Shift ends in the middle of run # 1482, radiator IN

    Runs:

    Run # radiator Evts Comments

    1464 IN 1M Probably bad
    1465 OUT 560k Probably bad, before Controlled Access
    1478 OUT 1M Maybe Good, but FPP Thresh=0
    1479 IN 1M Good
    1480 IN 700k Stopped due to beam trip
    1481 IN 300k
    1482 IN running